OpenStudy (anonymous):

Let f(x) = x^2-x. Find the average rate of change of f over the interval [0,t]. Simplify your answer.

OpenStudy (whpalmer4):

the average rate of change of a function \(f(x)\) over the interval \([a,b]\) is \[\frac{f(b)-f(a)}{b-a}\]which is just the slope of a line connecting the points \((a,f(a))\) and \((b,f(b))\)

OpenStudy (anonymous):

ok this is what I got average rate of change: f(t)-f(0)/t-0 = t^2-t - 0/t-0 = t^2-t/t

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

that's not simplified...

OpenStudy (whpalmer4):

also, if you are writing equations involving fractions without formatting them, you MUST use parentheses to indicate the proper grouping, such as (f(t)-f(0))/(t-0) = (t^2-t - 0)/(t-0) = (t^2-t)/t

OpenStudy (anonymous):

can you simplify (t^2-t) / 0? I was thinking that is the average rate of change

OpenStudy (anonymous):

(t^2-t) / t

OpenStudy (anonymous):

hmm, simplified would be just t^2

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

no...

OpenStudy (whpalmer4):

\[\frac{t^2-t}{t} = \frac{t^2}{t} - \frac{t}{t} = \]

OpenStudy (anonymous):

t^2/t

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

No, that's not correct. Are you telling me you don't know how to correctly reduce the fractions \[\frac{t^2}{t}\] and \[\frac{t}{t}\]?!?

OpenStudy (anonymous):

t^2 / t = t

OpenStudy (anonymous):

t^2-1 = t

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

I was not thinking about it when I put my answer... when you divide exponents, you subtract them

OpenStudy (anonymous):

I end up with t^2/t - t/t = t

OpenStudy (whpalmer4):

@jtryon still incorrect! \[\frac{t^2}{t} - \frac{t}{t} =\frac{t^2}{t^1} - \frac{t^1}{t^1} = t^{2-1} - t^{1-1} = t^1-t^0 = t - 1\]!!! I suggest you do some reviewing of exponents...